Find the integral surface passing through the initial data curve C: x0= -1, y0= s, z0= (s)1/2 of the equation (x+2)p +2yq =2z


1
Expert's answer
2021-01-01T14:54:51-0500

we have to find the integral surface of the equation :

"(x+2)p+2yq=2z..........(1)"

From observation we can see that it is "Lagrange's partial differential equation" of type

"Pp+Qq=R"

So, from equation"(1)" we get Lagrange's auxiliary equation

"\frac{dx}{x+2}=\frac {dy}{2y}=\frac{dz}{2z}\,.............(2)"

Now we take

"\frac{dx}{x+2}=\frac{dy}{2y}"

"\implies\intop\frac{dx}{x+2}=\intop\frac{dy}{2y}\\\implies\ln(x+2)=\frac{1}{2}\ln y+\ln c_1\\\implies\ln(x+2)-\ln(\sqrt{y})=ln c_1\\\implies\frac{x+2}{\sqrt{y}}=c_1............(3)"


we will take another equation

"\frac{dy}{2y}=\frac{dz}{2z}\\\implies\intop\frac{dz}{2z}=\intop\frac{dy}{2y}\\\implies\frac{1}{2}\ln z-\frac{1}{2}ln y=\ln\sqrt{c_2}\\\implies\sqrt{\frac{z}{y}}=\sqrt{c_2}\\\implies\frac{z}{y}=c_2..............(4)"

Now it is given that

"x_0=-1;y_0=s_0;z_0=\sqrt{s_0}\\"

putting the values of "x_0;y_0;z_0" in the equation (3) &(4) we get

"c_1=\sqrt{\frac{1}{s_0}}" & "c_2=" "\sqrt{\frac{1}{s_0}}"

now putting the values of C1& C2 we get

"\frac{x+2}{\sqrt{y}}=\sqrt{\frac{1}{s_0}}.......(5)" & "\frac{z}{y}=\sqrt{\frac{1}{s_0}}..........(6)"

now subtracting equation (5) to equation (6) we will get

"\frac{x+2}{\sqrt{y}}=" "\frac{z}{y}"

Hence

"\frac{x+2}{\sqrt{y}}=" "\frac{z}{y}" ; is the equation of the integral surface of the given equation.